Use substitution to solve for x1 and y1
y=x2
y=6x-5
i already figured out that x2=5 and y2=25. and yes i am sure those are correct

Found 2 solutions by Theo, solver91311:
Answer by Theo(13342)

you have y = x^2 and you have y = 6x - 5
since they are both equal to y, then they are both equal to each other and you get x^2 = 6x - 5
subtract 6x and add 5 to both sides of this equation to get x^2 - 6x + 5 = 0
this a quadratic equation whose factors are (x-5) * (x-1)
your equation becomes (x-5) * (x-1) = 0
since each of these factors can be set equal to 0, you get:
x-5 = 0 and x-1 = 0
solve for x in both these equations to get x = 5 and x = 1

Answer by solver91311(24713)



Since , you can write which is to say:



Factor and solve the quadratic for its two roots, one of which is, indeed, 5. Then substitute back into to find the two values of , one of which is, indeed, 25.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism